# Language Models can Self-Improve at State-Value Estimation for Better Search

### v1

@@ 4 Experiments
Reward and Demo Free Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 70.0 26.0 67.7 26.4
@@
4.1 Web Tasks
As mentioned in §1, it is particularly challenging and expensive to gather ground truth web task completion data (Zhang et al., 2024a) motivating self-improvement techniques like self-taught lookahead. To benchmark our self-taught lookahead method on web tasks, we utilize WebShop (Yao et al., 2022a) which consists of interactive web tasks involving searching for and purchasing an item that matches a short natural language specification. This benchmark is an ideal test bed to demonstrate the ability of our approach as, unlike other benchmarks (Zhou et al., 2023; Koh et al., 2024a), ground truth reward is provided for all tasks allowing a direct comparison between existing search and RL-based methods that use these rewards and self-taught lookahead which does not. Furthermore, WebShop also has a separate training set enabling the demonstration of self-improvement generalization and provides a textual representation of webpages that we leverage (see Appendix B.1 for an example).
Self-taught lookahead for web tasks.
Following the empirical advantages on agent tasks identified by Yu et al. (2024), we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) (Kocsis & Szepesvári, 2006) selection like Yu et al. (2024), instead of ground truth reward (Zhou et al., 2024). We perform self-taught lookahead with a gpt-3.5-turbo (Brown et al., 2020) policy to be consistent with previous work (Zhou et al., 2024; Shinn et al., 2024) as well as a gpt-4o (OpenAI, 2024) policy and fine-tune a llama-3.1-8b-instruct (Dubey et al., 2024) value function. STL is performed by rolling out 50 tasks from the WebShop training set resulting in close to 1500 training examples, which we find is sufficient for significant performance improvement. Additionally, while we perform data generation during self-taught lookahead with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the value of the policy’s proposed actions i.e. at each state $s_{i}$, we pick action $a_{i}$:
$a_{i}\leftarrow\max_{a\in A_{s_{i}}}\Big{\{}\text{Val}(V_{\phi}(x,s_{0},\cdots,s_{i})\Big{\}}$ (7)
As we will show in §5, greedy search enables a significant improvement in efficiency compared to prior methods.
Finally, we find that a single iteration of self-taught lookahead is sufficient to see significant improvement over using a base LLM-initialized value model and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead given the complexity of the environment (in §4.2 we show that this multi-step simulation is possible for simpler tasks). See Appendix B for further details about data generation, training, and further discussion about multiple iterations.
@@ 4.1 Web Tasks
Table 1 presents a comparison of the settings of the evaluated methods. We include baselines that use ground truth reward and a combination of search and reflection (Shinn et al., 2024; Zhou et al., 2024) by prompting closed-source LLMs, as well as imitation learning and RL methods that train smaller BERT (Devlin et al., 2019) and BART (Lewis et al., 2019) models on human demonstrations and ground truth reward (Yao et al., 2022a). We additionally compare to the current state-of-the-art approach AgentQ (Putta et al., 2024), which finetunes a xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization (Rafailov et al., 2023). Finally, we also include baselines in the Reward and Demo Free setting where only environment observations are accessible. These baselines include greedy search and MCTS [^3] We use LLM value as a proxy reward to guide UCT similar to the data generation phase of self-taught lookahead with a base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o). We also experiment with the recently released r1-distill-llama-8b (Guo et al., 2025), which allows a direct comparison between a general-purpose reasoning model and STL. We use the pass@$k$ (Chen et al., 2021) for methods that do not have access to ground truth reward at inference time [^4] Note the IL + RL method only has access to ground truth reward at train time..
@@
The WebShop average reward (Score) and success rate (SR) of all methods evaluated are presented in Table 2. We present results on both the full WebShop test set and on the mini test set of 50 tasks which was used by (Zhou et al., 2024) as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during self-taught lookahead. We find that self-taught lookahead leads to a greater than 17% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test (Berg-Kirkpatrick et al., 2012). Moreover, although it unsurprisingly underperforms compared to AgentQ, which is trained with environmental rewards, the STL model settings generally match or outperform tree search methods and settings where a closed-source LLM is used as a value model. Note that we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity which emphasizes the role of the value model over the policy (see Appendix B.2).
@@
We also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL, which uses the lookahead value, the textual representation of the next best action and the successor state, and the value rationale for the successor state, to other settings where only a subset of this information is used. The results of this ablation in Table 3 demonstrate that regressing solely on lookahead values and also incorporating state transitions from lookahead does improve performance relative to the base model. However, learning from the value rationale of the successor state, as done in STL, yields large improvements in success rate over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL (Gordon, 1999) and other LLM tree search works (Feng et al., 2023; Zhang et al., 2024a) which fine-tune an LLM value model on numerical values only.
Table 3: Ablation study on WebShop of the impact of fine-tuning with different combinations of information from lookahead namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
Fine-tuning Data Setup Score $\uparrow$ SR(%) $\uparrow$
llama-3.1-8b-instruct 70.4 30.0
    + lv 70.9 34.0
    + lv + tr 74.4 32.0
    + lv + tr + r (stl) 76.4 42.0
4.2 Math Reasoning
We also study the performance of self-taught lookahead on math reasoning through the Game-of-24 task where inputs consist of four integers between 1 and 13, and the goal is to construct a mathematical expression with the numbers to obtain 24. Performance on the Game-of-24 task has shown to benefit from tree search in previous works (Yao et al., 2024; Gandhi et al., 2024).
Figure 2: Breadth-first search performance on Game-of-24 task on sets of tasks both seen and unseen during the self-improvement.
Self-taught lookahead for math reasoning.
For the Game-of-24 task, we generate data using breadth-first search (BFS) rather than MCTS to be consistent with the original Tree-of-Thoughts approach. We use a gpt-4o policy and a llama-3.1-8b-instruct base value function in the first iteration. As described in §3, we replace this base value function with a trained model in each subsequent iteration. Self-taught lookahead is run for four iterations with 25 tasks used during each iteration. Note that in this case, we do not have explicit environment observations, but instead use the policy’s arithmetic while combining two numbers as a pseudo-observation. See Appendix C for further implementation details.
@@ 4.2 Math Reasoning
We compare the performance of value models learned via self-taught lookahead with BFS baselines that use the same gpt-4o policy but with varied value models. From the original setting proposed by Yao et al. (2024), one baseline uses the same LLM as the policy (gpt-4o) as the value model. Inspired by Chen et al. (2024), we also include a baseline with an algorithmic oracle evaluator. Specifically, the oracle runs a simple recursive algorithm to verify whether the current state (set of numbers) can be combined to reach $24$. Search performance with this oracle value gives us an upper bound on possible performance improvements due to improving the value function with a static policy.
@@
Figure 2 shows the performance of self-taught lookahead and the two baselines of two sets of 48 tasks. The first set includes tasks encountered during self-taught lookahead, comprising four sets of 12 tasks sampled from those observed in each iteration. The second set consists of more challenging tasks (determined by lower human solve percentages) that were not seen during self-taught lookahead.
On both evaluation sets, BFS with a STL llama-3.1-8b-instruct value model matches or outperforms BFS with a gpt-4o value model. However, self-taught lookahead’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing in later iterations. This phenomenon is likely due to the base value model generalizing better than self-improved models in the earlier iterations which have seen too few examples and have not received signal from more accurate LLM values of terminal states during lookahead and fine-tuning.
Figure 3: Compute and environmental efficiency during evaluation on WebShop with a gpt-3.5-turbo policy. Compute efficiency is measured in total (prompt and completion) tokens broken down by model type (closed and open source). Environmental efficiency is measured by the number of states expanded (webpages visited).
@@ 5 Efficiency Analysis
In this section, we compare the efficiency of self-taught lookahead with prior methods on the WebShop task. We study efficiency tradeoffs from two perspectives (1) model usage and costs and (2) environment usage. Additionally, in §5.3 we explore how performance changes when scaling the size of the value model trained during self-taught lookahead.
@@ 5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low during inference time is critical, especially for agents automating routine, repetitive tasks. Since self-taught lookahead can be used to improve an open-source LLM like llama-3.1-8b-instruct at value estimation, we can transfer computation from more resource-intensive and expensive closed-source models like gpt-4o to these open source models while maintaining performance. Figure 3 demonstrates this transfer, as self-taught lookahead uses more than 50% fewer closed source tokens as greedy search with a gpt-35-turbo or gpt-4o value model.
Figure 4: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between cost and average reward (top), environmental usage and average reward (middle), and cost and success rate (bottom).
We also compute the monetary cost of task performance using OpenAI’s555openai.com/api/pricing and Groq’s666groq.com/pricing - we only use Groq for a pricing estimate, as inference and training was run on in-house GPUs cost tables for closed-source models and llama-3.1-8b-instruct respectively. We plot these costs against the average WebShop reward and success rate in Figure 4 (top) and (bottom), respectively. We find that self-taught lookahead is Pareto optimal in both cases and is 37×37$\times$37 × cheaper than MCTS methods like LATS and 10×10$\times$10 × cheaper than performing greedy search with a gpt-4o value model.
@@ 5.2 Environmental Usage
Similar to sample efficiency during RL training, we propose that LLM agents should be environmentally efficient during inference. Specifically, in many cases, it may be crucial for an agent taking actions in physical or digital environments to be conservative in the number of states it visits while performing a task. In the case of digital web agents, taking many steps per task, for instance through exhaustive tree search, may put an unnecessary burden on web servers, especially as agents are deployed at scale. Additionally, allowing web agents to search the environment widely when equipped with personal information or the ability to make purchases may lead to unintended privacy disclosures or financial loss, respectively. Depending on the environment, taking a large number of actions may also directly lead to an unreasonable task completion time. Figure 3 and Figure 4 (middle) present the environmental efficiency measured by the count of expanded states or visited sites in the WebShop environment. Considering WebShop score, self-taught lookahead is Pareto optimal and requires expanding half as many states as MCTS-based methods like LATS. Furthermore, unlike LATS, self-taught lookahead does not require any irreversible actions (actually clicking Buy Now on a product page) required to obtain reward.
5.3 Self-Taught Lookahead Scaling Trends
Given that 8 billion parameter STL models can match the performance of a gpt-4o value model, is it possible to use even smaller models for STL while maintaining good performance? STL requires models to (1) provide generally consistent values out-of-the-box so that it is possible to compare successor states during data generation and (2) learn to generalize to unseen tasks and states, both of which may be challenging for smaller models. We explore this STL scaling trend on WebShop with $\leq$ 8 billion parameter models in the llama 3 family [^7] As there are no 1B and 3B models in the llama-3.1 family, we use llama-3.2-1b-instruct and llama-3.2-3b-instruct (Dubey et al., 2024) and $\leq$ 7 billion parameter models in the qwen-2.5-instruct family (Yang et al., 2024). The results presented in Figure 5 demonstrate that while performance does generally decrease with fewer parameters, smaller models like llama-3.2-3b-instruct and qwen-2.5-3b-instruct can match or even outperform the performance of gpt-4o in terms of average reward, respectively. This result indicates that these smaller models could be reasonably used for STL, increasing the feasibility of large-scale agent deployment to new domains.
Figure 5: STL scaling trends on WebShop for llama-3 and qwen-2.5 model families when using a gpt-3.5-turbo policy with performance measured by average reward (top) and success rate (bottom).
@@ 6 Related Work
Classical reinforcement learning.
STL is loosely inspired by the traditional fitted value iteration (FVI) (Gordon, 1999), which generalized value iteration (Bellman, 1957) beyond the tabular setting. In an iteration of FVI, target values are computed using the Bellman update and used to train a new value model from the previous model checkpoint using least squares regression. The iterated values in FVI are computed similarly to the lookahead values $y_{s_{k}}$ in §3.1, but with STL, no ground truth reward is assumed, the value model is non-Markovian, and actions are deterministic. Instead of learning directly from iterated values, with STL, they are concatenated with action-outcome rationales, and together, these sequences are used to fine-tune the LLM value model, training from scratch at each iteration rather than from the previous model checkpoint as in FVI.
@@
A variety of previous work has shown that LLMs can self-improve with iterative prompting techniques (Huang et al., 2022; Weng et al., 2022; Madaan et al., 2024) and have applied these methods to various domains, from agents (Kim et al., 2024) to privacy protection (Chen et al., 2023b). A separate line of work focuses on bootstrapping a small training dataset through a self-training process to improve either the reasoning policy model  (Gulcehre et al., 2023; Singh et al., 2023; Jung et al., 2023) or the verification or reward model (Hosseini et al., 2024) using synthetically generated data. While most self-training approaches utilize outcome-based reward models, other work (Aksitov et al., 2023; Zhang et al., 2024a) derive process-based rewards like STL to evaluate each step in the reasoning chain.
@@
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories (Yao et al., 2022a; Lai et al., 2024), synthetically generated trajectories (Chen et al., 2023a; Furuta et al., 2023; Zhang et al., 2024c; Liu et al., 2024b; Patel et al., 2024; Murty et al., 2024; Ou et al., 2024; Su et al., 2025), or a combination of both (Zeng et al., 2023; Zhang et al., 2024b). Other work has trained agents from tree search generated data, but these methods usually require ground truth reward (Gandhi et al., 2024; Putta et al., 2024; Zhang et al., 2024a). While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement (Patel et al., 2024) or require fine-tuning frontier models like gpt-4o to achieve generalization (Yu et al., 2024).

### v2

@@ 4.1 Web Tasks
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 70.0 26.0 67.7 26.4
@@
STL for web tasks.
Following the empirical advantages on agent tasks identified by yu2024exact , we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) kocsis2006bandit selection like yu2024exact , instead of ground truth reward zhou2023lats . We perform STL with a gpt-3.5-turbo gpt-35 policy to be consistent with previous work zhou2023lats ; shinn2024reflexion as well as a gpt-4o GPT-4o policy and fine-tune a llama-3.1-8b-instruct dubey2024llama value function. STL is performed by rolling out 50 tasks from the WebShop training set, resulting in close to 1500 training examples, which we find is sufficient for significant performance improvement. Also, while we perform data generation during STL with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the values of the policy’s proposed actions. Finally, we find that a single iteration of STL is sufficient to see significant improvement over using a base LLM-initialized value model, and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead, given the complexity of the environment. This conclusion has been corroborated by other LLM agent works gu2024your ; chae2024web . However, in §4.3 we show that this multi-step simulation is possible for simpler tasks. See Appendix D for further details about data generation, training, and simulation.
@@
Within Reward and Demo Learning approaches, we include IL and RL methods originally proposed in the WebShop work that train BERT devlin2019bert and BART lewis2019bart models on human demonstrations and ground truth reward yao2022webshop . We additionally compare to the current state-of-the-art approach AgentQ putta2024agent , which finetunes a larger xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization rafailov2023direct . We also include Reward-Guided Inference methods such as Reflexion shinn2024reflexion and LATS zhou2023lats , which work by prompting closed-source LLMs within a framework, e.g., MCTS, that is guided by ground truth reward. Finally, in the Reward and Demo Free setting, we include greedy search and MCTS [^2] We use LLM value as a proxy reward to guide UCT, like in the data generation phase of STL baselines with a ReACT yao2022react prompted base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o, r1-distill-llama-8b guo2025deepseek ). All value models are prompted with few-shot examples and are asked to provide reasoning before a numerical value. For clear comparisons, a branching factor of 5 is used for greedy and MCTS-based methods like LATS, i.e., LATS is strictly more computationally expensive than greedy methods. We use the pass@$3$ chen2021passk for methods that do not have access to reward at inference time.
@@
The WebShop average reward (Score) and success rate (SR) of all methods evaluated are presented in Table 1. We present results on both the full WebShop test set and on the mini test set of 50 tasks used by zhou2023lats , as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during STL. Within the Reward and Demo Free setting, we find that STL matches the performance of using a gpt-4o value model and leads to a greater than 7% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test berg2012empirical . Moreover, we find that STL even outperforms Reward-Guided Inference methods that have access to ground truth rewards. Lastly, we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity in our setup (see Appendix D.2); thus, we evaluate search methods with a single policy (either gpt-3.5-turbo or gpt-4o) in subsequent experiments.
Table 2: Ablation study on the impact of fine-tuning with different combinations of information from lookahead namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
Fine-tuning Data Setup Score $\uparrow$
llama-3.1-8b-instruct 70.0
+ lv 70.9
+ lv + tr 74.4
+ lv + tr + r (stl) 78.3
@@
In the Reward and Demo Free setting, we also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL to variants that use only subsets of the information derived from lookahead. As mentioned in §3, this information includes (1) the lookahead value, (2) the textual representation of the next-best action and its successor state, and (3) the value rationale for the successor state. The results of this ablation in Table 2 demonstrate that regressing solely on lookahead values and further incorporating state transitions from lookahead does improve performance relative to the base model. However, learning also from the value rationale of the successor state, as done in STL, yields additional performance gains over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL gordon1999fitted_value_iteration and other LLM tree search works feng2023alphazero ; zhang2024rest_mcts which fine-tune an LLM value model on numerical values only.
4.2 Multi-Hop Question Answering
We also investigate the efficacy of STL on applied reasoning for retrieval-based question-answering. We specifically utilize the HotpotQA yang-etal-2018-hotpotqa benchmark, consisting of multi-hop question-answering tasks that require retrieving and reasoning over multiple Wikipedia entries. We use the same setup as in §4.1 to generate data by rolling out with MCTS, but instead roll out 500 tasks from the training dataset since actions (search terms) proposed by the policy lack diversity compared to web tasks.
@@ 4.2 Multi-Hop Question Answering
In the Reward and Demo Learning setting, we evaluate on R1-Searcher song2025r1 , which performs RL on outcome-based answer correctness rewards on the multi-step retrieval QA task. As in §4.1, in the Reward-Guided Inference setting, we include Reflexion and LATS baselines, which have access to ground truth correctness during search. Among Reward and Demo Free methods, we compare STL to greedy search with closed-source models.
@@
Table 3 presents the answer match rate of all evaluated methods. Following feng2024agile , we prompt gpt-4o to help evaluate answer correctness. The prompt used is in Appendix E.3. This fuzzy matching correctness is exposed as the reward for Reflexion and LATS, but not to the greedy methods in the Reward and Demo Free setting. We evaluate on a set of 500 unseen questions and also a smaller set of 50 examples due to the high cost of LATS. STL on llama-3.1-8b-instruct outperforms a gpt-4o value model and the R1-Searcher method while approaching the performance of Reward-Guided Inference methods, which can verify predicted answer correctness during inference. Of the Reward and Demo Free approaches, STL is the only method that is statistically significant compared to the llama-3.1-8b-instruct baseline ($p<0.001$).
[…]
Setting Method Policy Value Match Rate $\uparrow$
Test Set (50) Test Set (500)
[…]
Reflexion gpt-3.5-turbo 70.0 66
[…]
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 60.0 56.4
gpt-3.5-turbo gpt-3.5-turbo 62.0 56.0
gpt-3.5-turbo gpt-4o 68.0 57.6
[…]
4.3 Math Puzzles
Figure 3: BFS Game-of-24 performance on tasks seen and unseen during STL.
Finally, we also study the performance of STL on the Game-of-24 task yao2024tree , where the goal is to construct a mathematical expression with 4 provided integers to obtain 24. The task was originally for Reward and Demo Free methods, so it serves as a good benchmark for STL.
STL for Game-of-24.
For this task, we generate data using breadth-first search (BFS) rather than MCTS to be consistent with the original Tree-of-Thoughts yao2024tree approach. We use a gpt-4o policy and a llama-3.1-8b-instruct base value function in the first iteration. As described in §3, we replace this base value function with a trained model in each subsequent iteration. STL is run for four iterations of 25 puzzles. For this task, we do not have explicit environment observations, but instead use the policy’s arithmetic to combine two numbers as a pseudo-observation. See Appendix F for further details.
Baselines.
We compare the performance of value models learned via STL with Reward and Demo Free BFS baselines that use the same gpt-4o policy. Specifically, we experiment with a gpt-4o value model and an algorithmic oracle evaluator chen2024search_useful . This oracle runs a recursive algorithm to verify whether the current state (set of numbers) can be combined to reach $24$. Search performance with this oracle is an upper bound on the performance improvement possible from improving the value function under a static policy.
Results and discussion.
Figure 3 shows the performance of evaluated methods on a set of 50 tasks seen during STL and a set of 50 more challenging (determined by lower human solve percentages), unseen tasks. On both sets, STL matches or outperforms a gpt-4o value model. However, STL’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing. This phenomenon may be due to the base value model generalizing better than self-improved models in earlier iterations, during which the model has not received signal from more accurate LLM values of terminal states during lookahead.
Figure 4: Compute and environmental efficiency during evaluation on WebShop with a gpt-3.5-turbo policy (left). Compute efficiency is measured in total (prompt and completion) tokens. Environmental efficiency is measured by the number of states expanded (webpages visited). The distribution of tokens (closed vs. open source models) used during search is also shown (right). Value models are specified in parentheses.
@@ 5 Efficiency Analysis
In this section, we compare the efficiency of STL with prior methods on WebShop. We study efficiency tradeoffs from two perspectives (1) model costs and (2) environment usage. Additionally, in §5.3 we explore how performance changes when scaling the size of the value model trained during STL.
@@ 5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low is critical, especially for agents automating routine, repetitive tasks. Since STL can be used to improve an open-source LLM like llama-3.1-8b-instruct, we can transfer computation from more expensive closed-source models like gpt-4o to open-source alternatives while maintaining performance. Figure 4 (right) demonstrates this transfer, as STL uses more than 50% fewer closed-source tokens than greedy search with a gpt-3.5-turbo or gpt-4o value model. We also compute the monetary cost of different methods, accounting for both inference costs and, for STL, those incurred during fine-tuning. See Appendix G for details about the cost calculation. We plot these costs against the average WebShop reward in Figure 5 (left), and find that STL is Pareto optimal, 23×23$\times$23 × cheaper than MCTS methods like LATS, and 5×5$\times$5 × cheaper than performing greedy search with a gpt-4o value model.
Figure 5: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between cost and average reward (left) and between environmental usage and average reward (right). Fine-tuning costs are included in the total STL cost.
@@ 5.2 Environmental Usage
It is often crucial for an agent taking actions in physical or digital environments to be environmentally efficient or conservative in the number of states it visits while performing a task. In the case of digital web agents, taking many steps per task through exhaustive tree search may put an unnecessary burden on web servers, especially as agents are deployed at scale. Additionally, allowing web agents to search widely when equipped with personal information or the ability to make purchases may lead to unintended privacy disclosures or financial loss, respectively. In some environments, taking many actions may also lead to unreasonable task completion times. Figure 4 (left) and Figure 5 (right) present the environmental efficiency measured by the count of expanded states or visited sites in the WebShop environment. Considering WebShop score, STL is Pareto optimal and requires expanding half as many states as MCTS-based methods like LATS. Moreover, unlike LATS, STL does not require irreversible actions (actually clicking Buy Now on a product page) required to obtain reward.
5.3 STL Scaling Trends
Figure 6: STL scaling trends on WebShop for llama-3 and qwen-2.5 families with a gpt-3.5-turbo policy.
As 8B STL models can match the performance of a gpt-4o value model, is it possible to use even smaller models for STL while maintaining good performance? STL requires models to (1) provide generally consistent values out-of-the-box so that it is possible to compare successor states during data generation and (2) learn to generalize to unseen tasks and states, both of which may be challenging for smaller models. We explore this STL scaling trend on WebShop with $\leq$ 8 billion parameter models in the llama 3 family [^3] As the llama-3.1 family lacks smaller models, we use 1B and 3B models from the llama-3.2 family. dubey2024llama and $\leq$ 7 billion parameter models in the qwen-2.5-instruct family yang2024qwen2 . The results presented in Figure 6 demonstrate that while performance does generally decrease with fewer parameters, the smaller 3B parameter models in both families do approach gpt-4o performance. This result suggests that using smaller models for STL is feasible, making large-scale agent deployment to new domains more practical.
@@ 6 Related Work
Classical RL.
STL is loosely inspired by fitted value iteration (FVI) gordon1999fitted_value_iteration , which generalized value iteration bellman1957dynamicprog to the tabular setting. In an iteration of FVI, target values are computed using the Bellman update and used to train a new value model from the previous model checkpoint using least squares regression. The iterated values in FVI are computed similarly to the lookahead values $y_{s_{k}}$ in §3.1, but with STL, no ground truth reward is assumed, the value model is non-Markovian, and actions are deterministic. Instead of learning directly from iterated values, with STL, they are concatenated with action-outcome rationales, and together, these sequences are used to fine-tune the LLM value model from scratch at each iteration rather than from the previous model checkpoint as in FVI.
@@
A variety of previous work has shown that LLMs can self-improve with iterative prompting techniques huang2022large ; weng2022large ; madaan2024self and have applied these methods to various domains, from agents kim2024language to privacy protection chen2023can . A separate line of work focuses on bootstrapping a small training dataset through a self-training process to improve either the reasoning policy model gulcehre2023rest ; singh2023rest_em ; jung2023impossible or the verification or reward model hosseini2024v using synthetically generated data. While most self-training approaches utilize outcome-based reward models, other work aksitov2023rest_react ; zhang2024rest_mcts derive process-based rewards like STL to evaluate each step in the reasoning chain.
@@
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories yao2022webshop ; lai2024autowebglm , synthetically generated trajectories chen2023fireact ; furuta2023multimodal ; zhang2024xlam ; liu2024apigen ; patel2024large ; murty2024nnetscape ; ou2024synatra ; su2025learn , or a combination of both zeng2023agenttuning ; zhang2024agentohana . Other work has trained agents from tree search generated data, but these methods usually require ground truth reward gandhi2024stream ; putta2024agent ; zhang2024rest_mcts . While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement patel2024large or require fine-tuning frontier models like gpt-4o to achieve generalization yu2024exact .